OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (5, -5) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

R=sqrt(5^2+5^2)=5*sqrt(2) Angle=tan^-1(-5/5)=-45 degrees OR another angle is -45+360=315 degrees